If 2.5m3 of a certain oil has a mass of 2turns. Find it's mass density
1
Expert's answer
2019-03-01T11:25:07-0500

Mass of sample (M)=2 tons=2000kg

Volume of the solution (V)=2.5m3

Mass density is defined as mass per unit volume

Mass density (d)=M/V=2000/2.5=800kg/m3